Evaluate h(-2/3) if h(x) =[[x-2]]
I get [[-2/3-2]]. Is the next step[[ -22/3]] or [[-1-2]]?
Please help
Thanks

Answer by AnlytcPhil(1810)

Evaluate h(-2/3) if h(x) =[[x-2]] 

Double brackets indicates the "floor function".
It means the largest integer that does 
not exceed the number between the double 
brackets:

h(x) =[[x-2]]

h(-2/3) = [[-2/3-2]]

h(-2/3) = [[-2/3 - 6/3]]

h(-2/3) = [[-8/3]]

h(-2/3) = [[-2.888888888888···]]

The largest integer that does not exceed
-2.888888888888···  is -3.

So h(-2/3) = -3

Edwin
AnlytcPhil@aol.com